OpenStudy (anonymous):

Solve the equation. Show your work. 16^3 = 4^x

OpenStudy (anonymous):

16^3 = 4^x 4^(2*3)=4^x 4^6=4^x x=6

OpenStudy (lgbasallote):

\[\Large 16^3 = 4^x\] \[\Large \implies (4^2)^3 = 4^x\] \[\Large \implies 4^6 = 4^x\] \[\Large \implies 6 = x\] does that help?

OpenStudy (anonymous):

kudos to you